OpenStudy (anonymous):

whether the ff each pair of equation are parallel, perpendicular or neither parallel nor perpendicular. 1.) x=-y+9 and y=-x-2

OpenStudy (whpalmer4):

You need to find the slope of each line. Easiest way to do that is to solve for \(y\), and then examine the coefficient of \(x\), which will be the slope. Parallel lines have equal slopes. If the slopes have a product of -1, the lines are perpendicular. Otherwise, they are neither parallel nor perpendicular unless one of them is of the form \(y = k\) and the other is of the form \(x = k\) where \(k\) is a constant (not necessarily the same value for each equation). A vertical line and a horizontal line are perpendicular, but a vertical line has an undefined slope, so you can't use the product of slopes = -1 test on it.

OpenStudy (anonymous):

give the equation example of no 1

OpenStudy (whpalmer4):

\[x = -y+9\]Add \(y\) to each side: \[x+y = -y + y + 9\]\[x+y = 9\]Subtract \(x\) from each side\[x+y-x = 9-x\]\[y = 9-x = -x + 9\] Slope is -1 because coefficient is -1. \(-x = -1x\)

OpenStudy (anonymous):

what is answer parallel

OpenStudy (whpalmer4):

I told you how to work it out, you need to do it yourself if you're going to learn anything.

OpenStudy (whpalmer4):

What is the slope of the other line?

OpenStudy (anonymous):

slope is -1

OpenStudy (anonymous):

tell the ff

OpenStudy (anonymous):

how to solve their coordinates

OpenStudy (whpalmer4):

Okay, so they have identical slopes. Identical slopes mean either parallel or the same line. Will the same point work in both equations? If so, that means they are the same line, otherwise parallel.

OpenStudy (anonymous):

in the 1st equation what is the slope 1